A string is 7.63 m long with a mass of 8.56 g. The string is held taut with a tension of 457 N applied to the string. A pulse is sent down the string. How long does it take the pulse to travel the 3.00 m of the string?


1
Expert's answer
2022-05-22T14:38:18-0400

The speed of wave

"v=\sqrt{T/(m/L)}\\
=\sqrt{457/(8.56*10^{-3}/7.63)}=638\:\rm m/s"

The required time

"t=d/v=3.00/638=4.70\:\rm s"